OpenStudy (anonymous):

Is there any rational numbers between 5/7 and 6/7 and why?

OpenStudy (anonymous):

I know there is, but I don't know why.

OpenStudy (anonymous):

Yes .... 5/7 can be rewritten as 10/14 6/7 can be rewritten as 12/14 and there is clearly a number between 10 and 12 .... but I do not know if this explains why?

OpenStudy (anonymous):

A rational number is defined as the ratio of two integers.

OpenStudy (anonymous):

So what meverett04 says is enough for me.

OpenStudy (across):

There's a theorem which states: For every \(x,y\in\mathbb{R}\) such that \(x<y\), there exists a rational number \(r\) such that \(x<r<y\).\[\]

OpenStudy (across):

Therefore, \(\mathbb{Q}\) is dense in \(\mathbb{R}\).\[\]

OpenStudy (anonymous):

Where can I find that kind of theorems @across?

OpenStudy (cwrw238):

i suppose there is one because difference between 6/7 and 5/7 can be exactly split into 2 12/14 - 10/ 14 / 2 = 1/14

OpenStudy (anonymous):

I mean books.

OpenStudy (anonymous):

can there be 11/14 between?

OpenStudy (across):

@No-data: Analysis books.

OpenStudy (anonymous):

yes there can be 11/14